### v1

# Spectral Hardening Reveals Afterglow Emergence in Long-Duration Fast X-ray Transients: A Case Study of GRB 250404A/EP250404a

Abstract
The prompt emission and afterglow phases of gamma-ray bursts (GRBs) have been extensively studied, yet the transition between these two phases remains inadequately characterized due to limited multiwavelength observational coverage. Among the recent growing samples of fast X-ray transients observed by Einstein Probe (EP), a subgroup of gamma-ray bursts are captured with long-duration X-ray emission, potentially containing featured evolution from prompt emission to the afterglow phase. In this Letter, we present a detailed analysis of GRB 250404A/EP250404a, a bright fast X-ray transient detected simultaneously by EP and Fermi/GBM in X-rays and gamma-rays. Its continuous X-ray emission reveals a long-duration tail, accompanied by distinct spectral evolution manifested by the spectral index $\alpha_{\rm X}$ with an initial softening, followed by an evident hardening, eventually reaching a plateau at the value of $\sim$ -2. Early optical and near-infrared observations enable broadband modeling with forward- and reverse-shock components, confirming that the X-ray hardening signals the emergence of the external-shock afterglow. From this spectral hardening we infer that the prompt phase in soft X-rays lasted $\sim 300\;\mathrm{s}$, which is more than three times longer than the gamma-ray $T_{90}$. This well-tracked soft-hard-flat spectral pattern provides a clear indication of afterglow emergence from the fading prompt emission and offers a practical criterion for identifying a distinct population of GRBs among fast X-ray transients, even when the detection of the gamma-ray counterpart or obvious temporal break is absent.
Transient sources; High energy astrophysics; X-ray transient sources; Gamma-ray bursts
1 Introduction
The afterglow phase of a gamma-ray burst (GRB) is conventionally defined as the stage that follows the end of the prompt emission (Zhang, 2018). From a theoretical perspective, the prompt emission is attributed to internal energy dissipation within the relativistic jet (Rees & Meszaros, 1994; Paczynski & Xu, 1994; Sari & Piran, 1997; Zhang et al., 2006), while the afterglow arises from the interaction between the ejecta and the circumburst medium (Paczynski & Rhoads, 1993; Mészáros & Rees, 1997). Observationally, the prompt emission typically exhibits rapid emission with erratic variability in the sub-MeV range, which is detected by GRB-triggering instruments (Zhang, 2014). In contrast, the afterglow is distinguished by broad-band spectra following a broken power-law shape and light curves displaying multi-segment broken power-law behavior, usually captured through multiwavelength follow-up observations (Sari et al., 1998). Extensive studies have explored the rich observational features of prompt emission, including its duration (Kouveliotou et al., 1993), spectral components (Zhang et al., 2011), and spectral evolution (Medvedev, 2006; Gao et al., 2021). Meanwhile, significant efforts have been devoted to understanding the afterglow phase, which is generally well-explained by synchrotron radiation originating from the external forward shock (FS) or the external reverse shock (RS) during the reverse shock crossing stage (Sari et al., 1998; Zhang & Kobayashi, 2005; Zhang et al., 2003).
However, the transition from the prompt emission phase to the afterglow phase is often not fully traced owing to a combination of instrumental, observational, and physical challenges. Instrumentally, the narrow field of view of most telescopes, except for GRB-triggering detectors, limits opportunities for obtaining simultaneous multiwavelength coverage during the prompt emission. Follow-up observations intended to capture the afterglow often begin with delays, resulting in incomplete coverage of the early emission period. Observationally, the curvature effect at the end of the prompt phase causes the emission to progressively soften and fade over time (Fenimore et al., 1996; Kumar & Panaitescu, 2000; Dermer, 2004), shifting it outside the energy range or below the sensitivity threshold of gamma-ray detectors. Physically, in the initial hours after the burst, the observed emission may be a complex superposition of multiple components (Zhang et al., 2006), including internal dissipation from late central engine activity, emission from the external RS, and from the external FS. As a result, it remains challenging to pinpoint when the afterglow emission from external shocks starts to appear and eventually dominates following the decay of the prompt emission.
The Einstein Probe (EP) Mission (Yuan et al., 2025), with its wide field of view, has detected dozens of fast X-ray transients, several of which exhibit prompt emission counterparts in gamma-rays (Yin et al., 2024; Liu et al., 2025; Jiang et al., 2025b), confirming their origin from the GRB internal energy dissipation of the relativistic jet. Notably, most of the EP-detected GRBs display soft X-ray tails that extend well beyond the duration of the gamma-ray emission, offering a valuable opportunity to study the spectral evolution during the late prompt emission and the transition to the afterglow emission. Crucially, the onboard trigger of EP/Wide-field X-ray Telescope (WXT) enables rapid localization with an accuracy of a few arcminutes, which is further refined to several arcseconds by the automatic follow-up observation from EP/Follow-up X-ray Telescope (FXT) within minutes. This capability facilitates timely multiwavelength follow-up observations by ground-based observatories during the early afterglow phase, concurrent with ongoing X-ray detection. Such synergy yields critical broadband insights into the distinct emission components. As a representative example, in this Letter, we present the EP-triggered detection of GRB 250404A/EP250404a featuring rich observations coverage X-ray and gamma-ray prompt emission, as well as optical and near-infrared follow-ups within few hours. A fast flux rise is observed in both X-ray and optical bands, following the spectral hardening in the X-ray emission. We performed the afterglow fitting on the multiwavelength data with a combined FS and RS model, and identified the very early afterglow emergence marked by clear spectral evolution in the X-ray band.
This Letter is organized as follows. Section 2 describes the observations and data reductions. Section 3 presents the analyses and results of the prompt emission and multiwavelength afterglow modeling. In Section 4, we summarize our findings and discuss different scenarios of the transition from the prompt emission to afterglow in the long-duration fast X-ray transients.
[…]
Observed Properties EP250404a
Redshift 1.88
Galactic $N_{\rm H}$ ($\rm cm^{-2}$) $6.00\times 10^{20}$
Intrinsic $N_{\rm H}$ ($\rm cm^{-2}$) $3.73_{-0.45}^{+0.35}\times 10^{22}$
$6.39_{-1.31}^{+1.61}\times 10^{21}$
Gamma-rays (10–1000 keV)
$T_{\rm 90}$ ($\rm s$) $90.43_{-0.37}^{+0.64}$
Spectral index $\alpha_{\gamma}$ ${-1.15}_{-0.02}^{+0.02}$
Peak energy (keV) ${55.34}_{-0.51}^{+0.64}$
Peak flux ($\rm erg\,cm^{-2}\,s^{-1}$) $8.62_{-0.28}^{+0.33}\times 10^{-7}$
Total fluence ($\rm erg\,cm^{-2}$) $3.37_{-0.01}^{+0.01}\times 10^{-5}$
Peak luminosity ($\rm erg\,s^{-1}$) $7.81_{-0.25}^{+0.30}\times 10^{51}$
Isotropic energy ($\rm erg$) $3.05_{-0.01}^{+0.01}\times 10^{53}$
X-rays (0.5–10.0 keV)
Duration ($\rm s$) $\sim 300$
Spectral index $\alpha_{\rm X}$ ${-2.59}_{-0.06}^{+0.05}$
Flux ($\rm erg\,cm^{-2}\,s^{-1}$) $9.61_{-0.23}^{+0.26}\times 10^{-9}$
Total fluence ($\rm erg\,cm^{-2}$) $1.20_{-0.01}^{+0.01}\times 10^{-6}$
Isotropic energy ($\rm erg$) $1.09_{-0.01}^{+0.01}\times 10^{52}$
  • * The parameters are derived in the time range 130–255 s.
[…]
2 Observations and Data Reduction
2.1 Einstein Probe Observations
[…]
The WXT cleaned event files and response files were generated following the standard data reduction pipelines implemented in the WXT Data Analysis Software (WXTDAS v2.10; Y. Liu et al., in preparation) and the calibration database (CALDB, v1.0; H.-Q. Cheng et al., in preparation). Source photons were extracted from a circular region with a radius of 999’9 ’, while background photons were extracted from an annular region with inner and outer radii of 181818’18 ’ and 363636’36 ’, respectively.
[…]
We note that WXT underwent a significant slew (indicated by the dashed curve in the first panel on the left of Figure 1) during the main emission phase. As a result, we caution that the WXT data presented in this work should be treated as qualitative references rather than precise measurements. The WXT light curve shown in Figure 1 is therefore intended to illustrate the temporal coverage provided by EP. Our quantitative analysis of the burst properties relies primarily on the FXT data.
2.2 Fermi/GBM Observations
The Fermi Gamma-ray Burst Monitor (GBM; Meegan et al., 2009) detected GRB 250404A (Mukherjee et al., 2025) at $T_{0}$, with a calculated location consistent with that of EP250404a (Fermi GBM Team, 2025). We retrieved the time-tagged event (TTE) data set covering the time range of GRB 250404A using the Python package heapy [^1] https://github.com/jyangch/heapy. From the 12 sodium iodide detectors onboard, we selected detectors n0, n1, and n9, which had the smallest viewing angles relative to the GRB source direction. In addition, we included the brightest bismuth germanium oxide detector, b0, to extend the energy coverage. Data reduction and analysis were performed with heapy, following the standard procedures described in Zhang et al. (2011) and Yang et al. (2022).
2.3 Ground-based Observations
We conducted ground-based follow-up observations in the optical and near-infrared bands using the Multi-channel Photometric Survey Telescope (Mephisto), the Tsinghua-Nanshan Optical Telescope (TNOT), the Sun Yat-sen University 80 cm telescope (SYSU 80cm), the Burst Observer and Optical Transient Exploring System (BOOTES)-4/María Eva Telescope (MET), and the Schmidt telescope. Additional imaging data were obtained from the the Alhambra Faint Object Spectrograph and Camera (ALFOSC) mounted on the 2.56 m Nordic Optical Telescope (NOT), located at the Roque de los Muchachos Observatory, La Palma, Spain, the Half-Meter telescope (HMT) located at Nanshan Observatory, Xinjiang, China, the JinShan 0.5 m (50D) and 1 m (100C) telescopes located at Altay Observatory, Xinjiang, China. Spectroscopic observations were carried out with the 2.4-meter telescope at the Yunnan Observatories (GMG-2.4m), covering the wavelength range of 3600–7460 Å. A summary of these observations is provided in Figure 3, Table LABEL:tab:photoag, Table LABEL:tab:photoclear and Figure 6.
2.3.1 Photometry
The optical counterpart of GRB 250404A/EP250404a was first detected by HMT at 14:24:59 UT on 2025 April 4, 313 s after the $T_{0}$. It was located at (J2000) R.A. = $08^{\rm hr}20^{\rm m}14.54^{\rm s}$, decl. = $+35^{\circ}31^{\prime}41.57^{\prime\prime}$ with an uncertainty of $0.5^{\prime\prime}$ (Jiang et al., 2025a). The observation lasted approximately 2.6 hours and consisted of a series of unfiltered exposures. We calibrated the photometry of the HMT data with Gaia DR3 $G$-band reference stars.
At $\sim$ 734 s after the $T_{0}$, Mephisto (Yang et al., 2024; Chen et al., 2024; Du et al., 2025) started the observation in the $u$, $v$, $g$, $r$, $i$ and $z$ bands. The optical afterglow was clearly detected at R.A. = 125.0607, decl. = 35.5282 (J2000). The $u$, $g$ and $i$ bands exhibited a rapid flux rise within the first $\sim$ 300 s of the observation, followed by a decay phase characterized by differing slopes between the early and late stages. Notably, the measurements in the blue bands of Mephisto deviate from the predictions of the standard afterglow model considering the correction for Galactic extinction along the line of sight (see Figure 4). This indicates the presence of additional extinction effects from the host galaxy or the surrounding medium.
We obtained a sequence of images in the $g,\,r,\,i,$ and $z$ bands using the 50D and 100C telescopes located at Altay Observatory. These observations were carried out between 0.8 and 4.6 hours after the $T_{0}$. Additionally, NOT/ALFOSC conducted follow-up observations in the same four filters approximately 8.3 hours after the $T_{0}$.
At $\sim$ $T_{0}$ + 1214 s, we also observed the field of the GRB 250404A/EP250404a with the 0.80m TNOT at Nanshan Station of Xinjiang Astronomical Observatory. A series of $g$, $r$ and $i$ band images were obtained, clearly detecting the optical afterglow at R.A. = 125.0606, decl. = 35.5282 (J2000). The fluxes derived from TNOT observations are in agreement with those measured by Mephisto in the corresponding optical bands.
In the $J$ band, Sun Yat-sen University 80cm infrared telescope started the observation at $\sim$ $T_{0}$ + 4021 s, obtaining 183 exposures of 20 s each. A counterpart was detected at the position of the optical afterglow in the stacked images.
Additional observations using clear filters have also captured the long-term decay features of the afterglow. Following the EP trigger, the 0.6m BOOTES-4/MET robotic telescope at Lijiang Astronomical Observatory automatically responded to this event at $\sim$ $T_{0}$ + 1203 s. A series of clear-filter images were collected, revealing a source consistent with the optical afterglow position. The magnitudes were measured using Gaia DR3 $G$-band as the reference. Furthermore, the 0.9 m Schmidt telescope at Xinglong Observatory, monitored the source over an extended period from $T_{0}$ + 1251 s to $T_{0}$ + 6379 s, capturing a series of unfiltered images. The magnitudes for these observations were calibrated using Gaia DR2 $G$-band as the reference.
2.3.2 Spectroscopy
The spectroscopic observation was conducted using the GMG-2.4m telescope with Grism #14, which provides a wavelength coverage of 3600-7460 Å(Wang et al., 2019; Xin et al., 2020). A single spectrum with an exposure time of 1800 s was obtained, beginning at $T_{0}$ + 3519 s. The spectrum was reduced using standard IRAF procedures (v2.18.1; Tody, 1986, 1993; Observatories & community, 2025). We identified multiple metal absorption features as shown in Figure 6, including Si IV at 1394 Å, Si IV at 1403 Å, Si II at 1527 Å, C IV at 1549 Å, Fe II at 1608 Å, Al II at 1671 Å, Al III at 1855 Å, Al III at 1863 Å, Zn II/Cr II at 2026 Å, Zn II/Cr II at 2062 Å, Fe II at 2344 Å, Fe II at 2374 Å and Fe II at 2383 Å. These features consistently indicate a redshift of z $\sim$ 1.88.
3 Prompt Emission to Afterglow transition
3.1 Early-Time Long-Duration Light Curve
In Figure 1, we present the light curve of GRB 250404A/EP250404a detected by Fermi/GBM with a bin size of 0.5 s in the energy range of 10–1000 keV. The accumulated counts curve is also shown, from which we extract the $T_{90}$ interval of ${90.43}_{-0.37}^{+0.64}$ s. The prompt emission phase of GRB 250404A is indicated by the $T_{90}$ and multi-pulse light curve with variability in this energy range.
We also display the scaled long-duration light curve detected by EP/WXT with a bin size of 3 s in the energy range of 0.5–4.0 keV. Continuous X-ray emission was affected by the slewing motion from $T_{0}$ + 15 s to 130 s and was disrupted at $T_{0}$ + 1384 s due to the end of the observation cycle. As a result, a reliable $T_{90}$ estimation in X-ray band for EP250404a could not be accurately derived. We also note that the second pulse of EP250404a appears broadened and extended compared to the gamma-ray emission, exhibiting a long tail with continuous emission that persists for thousands of seconds.
[…]
3.2 Spectral Evolution
We performed both time-integrated and time-resolved spectral fittings for Fermi/GBM and EP data using the Python package bayspec [^2] https://github.com/jyangch/bayspec (an upgraded version of MySpecFit), following the approach described in Yang et al. (2022) and Yang et al. (2023). bayspec is a Bayesian inference-based spectral fitting tool for multi-dimensional and multi-wavelength astrophysical data. The goodness of fit was evaluated by examining the reduced statistic STAT/dof, as described in Yin et al. (2024). The model comparison was conducted using the Bayesian information criterion (BIC) as defined by Schwarz (1978). The best-fitting model parameters, along with the corresponding statistics for each time slice, are provided in Table 3 and Table 4. Based on the spectral fitting results, we derived the spectral evolution and the spectral energy distributions (SEDs), which are illustrated in Figure 1.
Before analyzing the spectral evolution in detail, we firstly divided the early high-energy emissions detected by Fermi/GBM and EP into four time intervals based on data availability and performed time-integrated spectral fittings for each episode:
  • (a)
    EP/WXT and Fermi/GBM joint fit in $T_{0}$ + [-1, 15] s. Before the time interval when EP was slewing, there was a single overlapping time range of gamma-ray and X-ray emission, spanning $T_{0}$ + [-1, 15] s. We performed a joint fit using an absorbed cutoff power law (CPL) model, tbabs*ztbabs*cpl. In this model, tbabs and ztbabs represent the Tuebingen-Boulder ISM absorption model (Wilms et al., 2000), with parameters for the absorption column density, $N_{\rm H}$, and redshift. We adopted a Galactic absorption column density of $N_{\rm H,gal}\sim 6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$ and an intrinsic absorption column density of $N_{\rm H,int1}\sim 3.73\times 10^{22}\leavevmode\nobreak\ \rm{cm^{-2}}$, with the latter fixed to the value obtained from the time-integrated spectral fitting in $T_{0}$ + [130, 255] s using the EP/FXT spectrum (see below). The joint spectrum is well fitted, yielding a spectral index of ${-0.93}_{-0.04}^{+0.05}$ and a peak energy of $50.93_{-1.62}^{+1.55}$ keV. This time interval covers the first pulse of the prompt emission detected in gamma-rays, and the joint best-fit parameters represent an averaged view of the time-resolved spectral evolution during the first pulse (see below).
  • (b)
    Fermi/GBM independent fit in $T_{0}$ + [-4, 113] s. The prompt emission of GRB 250404A detected by Fermi/GBM expands from $\sim$ $T_{0}$4 s to +113 s. The time-integrated spectrum for this interval is well described by the CPL model, yielding a spectral index of ${-1.15}_{-0.02}^{+0.02}$ and a peak energy of ${55.34}_{-0.51}^{+0.64}$ keV. The isotropic energy derived from this episode is $3.05_{-0.01}^{+0.01}\times 10^{53}$ erg, placing GRB 250404A within the parameter space of Type II GRBs on the Amati relation (Amati et al., 2002).
  • (c)
    EP/FXT independent fit in $T_{0}$ + [130, 255] s and [255, 1384] s. The long-duration tail of EP250404a was detected by EP/FXT from $T_{0}$ + 130 s to + 1384 s. Considering the possibility of time-dependent absorption in the X-ray spectra, which is not uncommon in bright GRBs (Lazzati & Perna, 2002; Perna et al., 2003; Lazzati & Perna, 2003; Campana et al., 2007, 2021), we adopted different intrinsic absorption column densities for the time-resolved spectral fittings. However, the values of $N_{\rm H,int}$ for the time-resolved spectra are difficult to constrain accurately, with significant uncertainties. Thus, we determined the $N_{\rm H,int}$ from time-integrated spectra and fixed the value for time-resolved spectral fittings.
    We firstly performed a time-resolved spectral fitting test using the model tbabs*ztbabs*pl. In this model, the Galactic absorption column density $N_{\rm H,gal}$ was fixed at $6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$, while the intrinsic absorption column density $N_{\rm H,int}$ was treated as a free parameter. We divided the time-resolved spectra, ensuring each EP/FXT spectrum contained at least 300 total accumulating photon counts. Our analysis revealed significant evolution in the best-fit values of $N_{\rm H,int}$ between the fast-decaying phase ($\sim T_{0}$ + [130, 255] s) and the slow-decaying phase ($\sim T_{0}$ + [255, 1384] s) of the X-ray emission tail. Therefore, we adopted two distinct intrinsic absorption column densities determined from independent fit in $T_{0}$ + [130, 255] s and [255, 1384] s: $N_{\rm H,int1}$ $\sim$ $3.73\times 10^{22}$ $\rm cm^{-2}$ for the period before $T_{0}$ + 255 s and $N_{\rm H,int2}$ $\sim$ $6.39\times 10^{21}$ $\rm cm^{-2}$ for the period afterward.
    Accounting for the effect of an evolving $N_{\rm H,int}$ is non-trivial because the absorption in X-ray spectra is degenerate with the intrinsic spectral index. As shown in Figure 2, our treatment of the intrinsic absorption column density does not significantly affect the values of the spectral index, nor does it alter the overall trend of spectral index evolution.
Building on the time-integrated spectral fitting results, we performed time-resolved spectral fittings to derive the spectral evolution and SEDs, as shown in Figure 1. We briefly outline the time-resolved spectral fittings and analyses as follows:
  1. 1.
    The prompt emission in gamma-rays.
    We extracted the time-resolved spectra of Fermi/GBM in the time range of $T_{0}$ + [-4, 113] s to explore the spectral evolution of the prompt emission in gamma-rays. The time-resolved spectra were divided into two sets based on different criteria: one with a minimum of 30 average accumulated photon counts per channel and the other with a minimum of 120. The extracted spectra were fitted using the CPL model, and the spectral evolution was analyzed through the best-fit parameter. The resulting spectral evolution and SED plots are shown in Figure 1, where we observe that the peak energy, $E_{\rm p}$, exhibits a hard-to-soft transition during the first pulse, followed by an intensity tracking pattern (Golenetskii et al., 1983), with an overall declining trend. Additionally, the plots reveal that the spectral index, $\alpha_{\gamma}$, evolves with an overall softening trend, ranging from -0.51 to -1.58.
  2. 2.
    The long-duration tail in X-rays.
    The spectral evolution of the long-duration tail in X-rays was analyzed via time-resolved spectral fittings with fixed intrinsic absorption column density $N_{\rm H,int1}$ and $N_{\rm H,int2}$ determined from time-integrated spectral fittings. We employed two spectral models, an absorbed powerlaw (PL) tbabs*ztbabs*pl and an absorbed smooth broken powerlaw (SBPL) tbabs*ztbabs*sbpl, for the time-resolved spectral fittings of EP/FXT spectra. Most spectra were better described by tbabs*ztbabs*pl. However, for the time-resolved spectra in the intervals $T_{0}$ + [130, 154] s and $T_{0}$ + [290.5, 462.0] s, both models provided comparable fits ($\Delta\rm{BIC}\textless 5$). We identified the spectral evolution of the peak energy passing through EP/FXT 0.5–10.0 keV band within $T_{0}$ + [130, 154] s, as revealed by the SBPL model, which aligns with the softening trend observed in Fermi/GBM spectral fittings in $T_{0}$ + [-4, 113] s. Furthermore, the spectral indices $\alpha_{\rm X}$ demonstrate significant hardening after $T_{0}$ + 255 s, eventually reaching a plateau at $\sim-2$. Notably, the first two time-resolved spectra after $T_{0}$ + 255 s clearly display a transition, as seen in the SBPL model in the SED plot. This suggests the emergence of a second spectral component, distinct from the softening and fading prompt emission.
Lastly, joint fittings were performed using EP/FXT-A and EP/FXT-B spectra for the time intervals $T_{0}$ + [44412, 47458] s and $T_{0}$ + [114594, 122326] s. We incorporated a calibration constant for EP/WXT-A to account for the systematic uncertainty between the two detectors. The X-ray flux densities were then derived for the afterglow fitting.
Figure 3: Multiwavelength observations of GRB 250404A/EP250404a and afterglow modeling with best-fit parameters from the FS+RS model with $u$, $v$, and $g$ band correction factors. The multiwavelength data utilized in the afterglow fitting are marked in solid points, while the optical observations on clear filters that are not included in the afterglow fitting are marked in hollow points. The optical and near-infrared data have been corrected for Galactic extinction, which is $E(B-V)=0.073$ mag (Schlafly & Finkbeiner, 2011). We assume a total to selective extinction ratio of $R_{V}=3.1$ according to the extinction law from Fitzpatrick (1999). The inverted triangle points signify upper limits. The best-fit FS+RS, FS and RS models are shown with solid, dashed and dotted lines, respectively. The dark blue dashed vertical lines corresponds to the same time interval where the spectra start to harden and reach a plateau.
Figure 4: X-ray/optical/near-infrared afterglow specific flux density spectra of GRB 250404A/EP250404a in different time intervals. The best-fit afterglow model is indicated by dashed gray lines. The red curves correspond to the partial host galaxy extinction curve derived from the best-fit correction factors from the afterglow modeling. For comparison, the yellow, blue and green curves rshow the average extinction laws of the MW, SMC, LMC, respectively, assuming a reddening value of $E(B-V)=0.1$ mag.
3.3 Confirming Afterglow Emergence via Multiwavelength Fitting
Corresponding to the spectral hardening phase starting at $\sim$ $T_{0}$ + 255 s, the flux density at 1 keV initially exhibits a brief decay, followed by a sharp rise peaking at $\sim$ $T_{0}$ + 1000 s. This rising flux is also observed in the $u$, $g$ and $i$ band of Mephisto early-time optical data. Subsequently, the flux density shows a steeper slope before $\sim$ $T_{0}$ + 5000 s and a shallower decay at later times. These behaviors indicate the presence of a RS component emerging around $T_{0}$ + 1000 s, superimposed on the FS emission.
We fitted the multiwavelength data with both FS and FS+RS models using PyFRS [^3] https://github.com/leiwh/PyFRS (Zhang, 2018; Gao et al., 2013; Lei et al., 2016; Zhu et al., 2023). Given the potential host galaxy extinction affecting the blue bands, we adopted two approaches: (1) fitting the afterglow without the heavily affected $u$ and $v$ bands, and (2) fitting the afterglow with extinction correction factors applied to the $u$, $v$, and $g$ bands.
The fitting of the multiwavelength data is implemented by the Bayesian computation python package PyMultinest (Buchner et al., 2014) with the log-likelihood function written as
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{n}\left\{\frac{(O_{i}-P_{i})^{2}}{\sigma_{i}^{2}+v^{2}}+\ln\left[2\pi(\sigma_{i}^{2}+v^{2})\right]\right\},$ (1)
where $O_{i}$, $P_{i}$, and $\sigma_{i}$ stand for the $i$th of $n$ observed magnitudes, model-predicted magnitudes, and the uncertainties of observed magnitudes, respectively. An additional variance parameter $v$ is introduced to as a scatter term, which accounts for additional uncertainty in models and data. For upper limits, a one-sided Gaussian penalty term is applied. To measure the goodness of fit between the model and the observed data, we utilize $\chi^{2}$ as the statistical metric. All model parameters are allowed to vary, with broad but physically reasonable priors, enabling a comprehensive search for the best-fit solution.
The best-fitting model parameters and their corresponding statistics for the four fitting scenarios are presented in Table 2. Among these, only the FS+RS model with extinction correction factors accurately reproduces the multiwavelength data. The corner plot of the posterior probability distributions of the parameters are presented in Figure 7. Furthermore, the derived extinction correction factors are consistent with the specific flux density spectra shown in Figure 4 in different time slices, which formed a steeper curve compared to the average extinction laws of Milky Way (MW), Small Magellanic Clouds (SMC) and Large Magellanic Clouds (LMC), assuming a reddening value of $E(B-V)=0.1$.
These results confirm that the second spectral component emerging after $T_{0}$ + 255 s corresponds to the FS+RS afterglow emission of GRB 250404A/EP250404a.
[…]
(erg) () $\rm(cm^{-3})$
FS+RS ${55.52}_{-0.78}^{+0.23}$ ${2.38}_{-0.23}^{+0.06}$ ${6.04}_{-2.11}^{+2.74}$ ${-1.08}_{-0.46}^{+1.30}$ ${2.86}_{-0.04}^{+0.01}$ ${-2.10}_{-0.04}^{+0.66}$ ${-4.11}_{-1.25}^{+0.39}$ ${2.64}_{-0.01}^{+0.05}$
FS ${53.35}_{-0.02}^{+0.54}$ ${2.89}_{-0.24}^{+0.06}$ ${9.67}_{-0.23}^{+0.27}$ ${3.52}_{-0.15}^{+0.65}$ ${2.81}_{-0.01}^{+0.02}$ ${-0.28}_{-0.55}^{+0.00}$ ${-6.32}_{-0.54}^{+0.24}$
FS+RS ${53.07}_{-0.01}^{+0.21}$ ${1.75}_{-0.04}^{+0.07}$ ${6.96}_{-0.56}^{+1.06}$ ${2.43}_{-0.14}^{+0.52}$ ${2.68}_{-0.01}^{+0.03}$ ${-0.52}_{-0.07}^{+0.00}$ ${-5.15}_{-0.44}^{+0.05}$ ${2.26}_{-0.09}^{+0.04}$
FS ${53.71}_{-0.21}^{+0.54}$ ${2.91}_{-0.31}^{+0.03}$ ${9.81}_{-0.74}^{+0.09}$ ${3.81}_{-0.87}^{+0.25}$ ${2.82}_{-0.01}^{+0.04}$ ${-0.51}_{-0.58}^{+0.05}$ ${-6.81}_{-0.04}^{+0.84}$
[…]
FS+RS ${-0.60}_{-0.38}^{+0.34}$ ${-5.45}_{-0.32}^{+0.73}$ ${-0.84}_{-0.02}^{+0.02}$ ${-1.14}_{-0.01}^{+0.06}$ ${-0.57}_{-0.02}^{+0.02}$ ${-0.20}_{-0.02}^{+0.00}$ 353.41/336 -178.19
FS - - ${-0.64}_{-0.01}^{+0.02}$ ${-1.13}_{-0.02}^{+0.07}$ ${-0.55}_{-0.03}^{+0.03}$ ${-0.22}_{-0.02}^{+0.01}$ 345.55/339 77.11
FS+RS ${-0.66}_{-0.11}^{+0.08}$ ${-4.01}_{-0.44}^{+0.03}$ ${-0.57}_{-0.00}^{+0.02}$ 322.69/314 160.20
FS - - ${-0.51}_{-0.01}^{+0.03}$ 330.84/317 228.74
  • * Three constants are added to the model for the host galaxy extinction corrections on $u$, $v$ and $g$ bands.
4 Summary and Discussion
In this Letter, we present a detailed analysis of the long-duration emission from the fast X-ray transient EP250404a, whose gamma-ray counterpart was simultaneously detected as GRB 250404A (see Table 1). The wide-field detector EP/WXT captured the prompt emission phase of GRB 250404A, while the rapid automatic follow-up of EP/FXT recorded a continuous emission tail lasting for thousands of seconds after the trigger. We conducted a comprehensive spectral analysis of the long-duration emission in gamma-ray and X-ray bands, revealing a complete spectral evolution from the prompt emission to the afterglow. The spectral peak energy decreased from above 100 keV to below 0.5 keV within the first $\sim$ 154 s from the trigger. The spectral index $\alpha_{\rm X}$ showed a softening trend during the decaying tail before $T_{0}$ + 255 s, followed by a clear hardening, eventually reaching a plateau at $\sim$ -2. This spectral hardening coincides with a sharp flux rise at 1 keV and in the optical bands, peaking at $\sim$ $T_{0}$ + 1000 s. The subsequent decay in flux density is characterized by distinct slopes in the early and late stages. Our multiwavelength data are best fitted by a FS+RS afterglow model with extinction correction factors in the blue bands, confirming that the emerging second component at $T_{0}$ + 255 s, indicated by the spectral hardening, is the afterglow emission. These findings also provide an estimation of the prompt emission phase duration in X-rays, lasting $\sim$ 300 s, after which the X-ray emission is dominated by the afterglow.
Interestingly, spectral hardening has become less uncommon in the detection of long-duration fast X-ray transients by EP, thanks to its automatic rapid follow-up capability, which enables the identification of an increasing number of events exhibiting these features. This phenomenon, characterized by the evolution of peak energy $E_{\rm p}$, softening and hardening of the spectral index $\alpha_{\rm X}$, and the appearance of spectral breaks in the SEDs, may signify a universal transition from the prompt emission phase to the afterglow phase. Benefiting from simultaneous gamma-ray observations and extensive multiwavelength follow-ups in the optical and near-infrared bands, we were able to conduct an in-depth analysis of the spectral evolution for this event. In this context, GRB 250404A/EP250404a may serve as a representative case of a specific subgroup of fast X-ray transients, distinguished by a continuous decay phase exhibiting both spectral softening and subsequent hardening. The physical mechanisms explored in this study could potentially be applicable to other similar events.
These findings suggest that the transition between the prompt emission and afterglow phases can occur in various forms, depending on the relative strengths and durations of the two components. We briefly summarize three typical scenarios (as illustrated in Figure 5) for the transition between prompt emission and afterglow in GRBs that are observed in long-duration fast X-ray transients:
  • Case I: The prompt emission and afterglow components are superimposed, with the prompt emission exhibiting a harder spectral index than the afterglow at the point when the afterglow begins to dominate. In this scenario, a gradual softening of the spectral index is expected. However, the spectral overlap makes it difficult to confidently identify the afterglow component or confirm the GRB-dominated nature of the fast X-ray transient.
  • Case II: The prompt and afterglow emissions are superimposed and of comparable intensity. The presence of the afterglow broadens the fast decaying profile of the prompt emission. In this case, the spectral index of the prompt emission softens beyond the value of the afterglow (we take $\sim$ $-2$ for the demonstration) before the afterglow becomes dominant. This results in a characteristic soft–hard–flat pattern in the spectral index evolution, accompanied by a smoothly decaying light curve. The spectral evolution in this case is particularly informative for identifying the emergence of the afterglow. The prompt emission and afterglow components can be disentangled by fitting the time-resolved spectra during the spectral hardening phase with two power-law models, allowing for the determination of their respective fluxes. While the total decay light curve is featureless, the decay slopes of the separated components align with standard GRB expectations: a steep decay for the prompt emission and a shallower, typical decay for the afterglow.
    Figure 5: The spectral index evolutions and light curves for three scenarios of the transition from the prompt emission to the afterglow in the long-duration fast X-ray transients. The purple, blue and yellow dots represent the spectral indices of the total, prompt and afterglow emission spectra. The colored curves demonstrate the flux evolution in linear and logarithmic scale. Upper panel: Case I. Middle panel: Case II. Lower panel: Case III.
  • Case III: The afterglow emerges at later times, after the prompt emission has already softened and faded. Spectral hardening is observed in the long-duration decaying tail of the continuous X-ray emission, with the flux exhibiting two distinct phases: an initial steep decay followed by a transition to a normal decay slope. This behavior is observed in this work.
In summary, this study identifies a distinct spectral evolution pattern in fast X-ray transients, where the spectral index $\alpha_{\rm X}$ transitions from softening to hardening, eventually reaching a plateau. This pattern directly signals the emergence of the afterglow following the prompt emission phase. Importantly, it provides a reliable criterion for distinguishing a specific subgroup of GRBs among fast X-ray transients, even in the absence of gamma-ray counterparts or clear temporal features in the X-ray flux.
We are grateful to Xinglong-2.16m telescope team for the helpful discussions on the redshift determination. We also thank the support of the YNAO staff at Lijiang Observatory. We acknowledge the support by the National Key Research and Development Programs of China (2022YFF0711404, 2022SKA0130102, and 2021YFA0718500), the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 11833003, U2038105, U1831135, 12121003, 12393811, and 13001106), the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-B11, and the Fundamental Research Funds for the Central Universities. This work is based on data obtained with the Einstein Probe, a space mission supported by the Strategic Priority Program on Space Science of the Chinese Academy of Sciences, in collaboration with ESA, MPE and CNES (Grant No. XDA15310000, No. XDA15052100). Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for WorldClass University” and “Yunnan University Development Plan for World-Class Astronomy Discipline”. The authors from YNU acknowledge support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” - the “Innovation Team” (202105AE160021) and the “Top Team” (202305AT350002), all funded by the “Yunnan Revitalization Talent Support Program”. This work is also supported by the National Key Research and Development Program of China (2024YFA1611603) and the “Yunnan Provincial Key Laboratory of Survey Science” with project No. 202449CE340002. Y. Fang is supported by the Yunnan Fundamental Research Projects (202301AU070006). X. Wang is supported by NSFC (12288102, 12033003) and the New Cornerstone Science Foundation through the XPLORER PRIZE. Members of TNOT acknowledges financial support from the Natural Science Foundation of Xinjiang Uygur Autonomous Region under No. 2024D01D32 and the Tianshan Talent Training Program grant 2023TSYCLJ0053 and 2023TSYCCX0101, the Central Guidance for Local Science and Technology Development Fund under No. ZYYD2025QY27, the National Natural Science Foundation of China NSFC 12433007.
The SYSU 80cm infrared telescope is operated and managed by the Department of Astronomy, Sun Yat-sen University. J.M. has the financial support of the National Key R&D Program of China (2023YFE0101200), Natural Science Foundation of China 12393813, and the Yunnan Revitalization Talent Support Program (YunLing Scholar Project). AJCT acknowledges support from the Spanish Ministry project PID2023-151905OB-I00 and Junta de Andalucía grant P20_010168 and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. DBM is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140.
Appendix A Spectral Fitting Results
[…]
t1 (s) t2 (s) CPL model
[…]
(keV) ($\rm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ keV^{-1}}$)
-4.00 113.00 ${-1.15}_{-0.02}^{+0.02}$ ${55.34}_{-0.51}^{+0.64}$ ${-1.54}_{-0.02}^{+0.01}$ 749.19/219 765.4
-1.00 15.00 ${-0.93}_{-0.04}^{+0.05}$ ${50.93}_{-1.62}^{+1.55}$ ${-1.45}_{-0.04}^{+0.05}$ 218.15/187 233.89
15.00 23.00 ${-0.97}_{-0.04}^{+0.04}$ ${76.38}_{-1.57}^{+1.78}$ ${-1.16}_{-0.03}^{+0.03}$ 277.31/213 293.44
23.00 31.00 ${-0.87}_{-0.06}^{+0.05}$ ${54.20}_{-1.11}^{+1.13}$ ${-1.11}_{-0.06}^{+0.04}$ 225.80/186 241.52
31.00 46.00 ${-0.73}_{-0.06}^{+0.05}$ ${50.58}_{-0.81}^{+0.94}$ ${-1.07}_{-0.05}^{+0.05}$ 248.63/182 264.29
46.00 54.00 ${-1.24}_{-0.06}^{+0.05}$ ${73.28}_{-2.32}^{+3.28}$ ${-1.50}_{-0.05}^{+0.04}$ 228.99/203 244.98
54.00 71.00 ${-1.39}_{-0.04}^{+0.06}$